# supernovae: individual (SN 1181) — supernovae: general — white dwarfs — accretion — hydrodynamics

### v1

@@ 1 Introduction
Among historical SNRs, the most recent identification concerns the SN recorded in 1181. 2021ApJ...918L..33R reported the infrared nebula Pa 30 (also IRAS 00500+6713) roughly in the location of the historical SN, and from its expansion velocity ($\sim 1000~\mathrm{km~s^{-1}}$) and size ($\sim 1$ pc) inferred an age of $\sim 1000$ yr, making Pa 30 the leading candidate counterpart to the SN 1181 remnant. At its center lies a peculiar object that drives an optically thick wind with a terminal velocity $\sim 1.5\times 10^{4}~\mathrm{km~s^{-1}}$ and a mass-loss rate of $\sim 10^{-6}~M_{\odot}~\mathrm{yr^{-1}}$, as inferred from optical spectra (2019Natur.569..684G; Lykou2022). The photospheric composition, combined with the wind velocity, suggests that an oxygen–neon (ONe) white dwarf (WD) (2019Natur.569..684G; 2019ApJ...887...39K; Lykou2022) exists inside.
@@
Type Iax SNe are a subclass of thermonuclear SNe characterized by low peak luminosities (absolute magnitudes $\sim-13$ to 18-18), slow expansion velocities ($\sim 6000$$7000~\mathrm{km~s^{-1}}$ at maximum light), and the absence of the near-infrared secondary maximum seen in normal SNe Ia (see 2017hsn..book..375J, for a review). They are often categorized into “bright” and “faint” groups (see Fig. 1). The leading theoretical model invokes a pure-deflagration in a WD: the flame does not transition to a detonation and propagates subsonically, yielding lower explosion energies and typically a bound remnant. Pure-deflagration models of CO WDs reproduce many observed properties of the bright Iax events (2012ApJ...761L..23J; 2013MNRAS.429.2287K; 2014MNRAS.438.1762F; 2022A&A...658A.179L), while hybrid CONe WDs can account for bright events and some faint ones (2015MNRAS.450.3045K; 2023ApJ...959..112F). In this framework, the deflagration is triggered by accretion of He-rich material from a He star. HST pre-explosion imaging of the type Iax SN 2012Z revealed a He-rich companion candidate (2014Natur.512...54M), and the source persists after the explosion (2022ApJ...925..138M), consistent with being a surviving donor. However, reproducing the faintest SNe Iax within pure-deflagration models often requires very small ejecta masses, in tension with SN 2021fcg—the faintest Iax yet observed—whose inferred ejecta mass is comparable to that of brighter events (2021ApJ...921L...6K).
@@
An alternative channel is an ONe + CO WD merger in which only the CO component detonates, leaving the ONe WD as the remnant (2018ApJ...869..140K); this pathway generally does not reproduce light curves and spectra of SNe Iax as well as pure-deflagration models. Moreover, recently identified intermediate-luminosity events such as SN 2019muj (2021MNRAS.501.1078B; 2021PASJ...73.1295K; 2022ApJ...941...15M), SN 2022xlp (2025arXiv250907717B), SN 2024pxl (2025arXiv250502943S; 2025ApJ...988..209H; 2025ApJ...989L..33K) and SN 2025qe (2025arXiv250602118M) show features overlapping with those of both the bright and faint groups, suggesting a continuous character rather than two discrete subclasses. Taken together, these considerations keep the progenitor systems of SNe Iax under active discussion. Recent work further argues that in single-degenerate systems a deflagration–detonation transition may often fail, favoring Iax-like, low-energy outcomes over normal SNe Ia (2025arXiv250716907M). Clarifying the progenitors of SNe Iax is therefore central to the broader SN Ia progenitor problem.
@@
Figure 1: Decline rate $\Delta m_{15}$ versus peak absolute magnitude $M$ for normal SNe Ia and for the “bright” and “faint” subclasses of SNe Iax (reproduced from Fig. 11 of 2022MNRAS.511.2708S; filled/open symbols denote rr/$R$). The red region marks SN 1181; its values are not measured in $r/R$ but inferred from historical visual (naked-eye) estimates: a peak $M\simeq-12.5$ to 14-14 mag (2021ApJ...918L..33R). The SN 1181 region is placed on the $r/R$ plane for comparison only.
@@
Several recently studied SNe Iax remnants show optically thick winds (e.g., 2016MNRAS.461..433F; 2023ApJ...951...67C; 2025arXiv250401063S), as seen in Pa 30, likely powered by delayed radioactive decays of 56Ni and 56Co (2017ApJ...834..180S). Their characteristic velocities ($\sim 500$$1000~\mathrm{km~s^{-1}}$) are, however, far below the extreme wind inferred for the central WD in Pa 30. To account for this exceptionally fast wind, 2019ApJ...887...39K proposed a model involving a rapidly rotating, strongly magnetized WD that accelerates a carbon-ash outflow where the main energy source is shell carbon burning; the evolutionary pathway to such a configuration has not been systematically explored.
@@ 2 Methods
We investigate conditions for the ignition of fallback matter accreted onto the central WD inside Pa 30 to find the parameter space that allows for the delayed carbon ignition on timescales of $\sim 800$ yr. As discussed in Sec 1, the pure-deflagration model remains the leading progenitor scenario for SNe Iax, while merger channels involving an ONe WD and CO WD have also been proposed. Because simulations across these scenarios indicate that SN Iax ejecta are CO-rich (e.g., 2014MNRAS.438.1762F; 2018ApJ...869..140K), we restrict our analysis to accretion of matter containing only CO for simplicity. In pure-deflagration scenarios, the bound remnant at the center may be either CO or ONe WD, whereas in the WD merger channel the central remnant is generally an ONe WD. For Pa 30, prior estimates place the WD mass at $\sim 1.1$$1.3~M_{\odot}$ (e.g., 2019ApJ...887...39K; Lykou2022), which also prefers an ONe WD according to the stellar evolution calculations.
@@ 4.1 Implication on SN 1181 Progenitor system
If SN 1181 is a pure-deflagration SN Iax, a surviving He star companion is expected, however, none has been identified within Pa 30. Two possibilities remain at least. First, if the WD is enshrouded by an optically thick wind with the photosphere at $\sim 10^{10}~\mathrm{cm}$, the He star could lie within it and be obscured. Second, the He star companion may lie outside the photosphere but be intrinsically faint—particularly at UV wavelengths—and has therefore escaped detection in existing observations. The two cases are examined in the following subsections.

### v2

@@ 1 Introduction
Among historical SNRs, the most recent identification concerns the SN recorded in 1181. 2021ApJ...918L..33R reported the infrared nebula Pa 30 (also IRAS 00500+6713) roughly in the location of the historical SN, and from its expansion velocity ($\sim 1000~\mathrm{km~s^{-1}}$) and size ($\sim 1$ pc) inferred an age of $\sim 1000$ yr, consistent with the estimate by X-ray observation (Oskinova_et_al_20), making Pa 30 the leading candidate counterpart to the SN 1181 remnant. At its center lies a peculiar object that drives an optically thick wind with a terminal velocity $\sim 1.5\times 10^{4}~\mathrm{km~s^{-1}}$ and a mass-loss rate of $\sim 10^{-6}~M_{\odot}~\mathrm{yr^{-1}}$, as inferred from optical spectra (2019Natur.569..684G; Lykou2022). The photospheric composition, combined with the wind velocity, suggests that an oxygen–neon (ONe) white dwarf (WD) (2019Natur.569..684G; 2019ApJ...887...39K; Oskinova_et_al_20; Lykou2022) exists inside.
@@
Type Iax SNe are a subclass of thermonuclear SNe characterized by low peak luminosities (absolute magnitudes $\sim-13$ to 18-18), slow expansion velocities ($\sim 6000$$7000~\mathrm{km~s^{-1}}$ at maximum light), and the absence of the near-infrared secondary maximum seen in normal SNe Ia (see 2017hsn..book..375J, for a review). They are often categorized into “bright” and “faint” groups (see Fig. 1). The leading theoretical model invokes a pure-deflagration in a WD: the flame does not transition to a detonation and propagates subsonically, yielding lower explosion energies and typically a bound remnant. Pure-deflagration models of CO WDs reproduce many observed properties of the bright Iax events (2012ApJ...761L..23J; 2013MNRAS.429.2287K; 2014MNRAS.438.1762F; 2022A&A...658A.179L), while hybrid CONe WDs can account for bright events and some faint ones (2015MNRAS.450.3045K; 2023ApJ...959..112F). In this framework, the deflagration is triggered by accretion of He-rich material from a He star. HST pre-explosion imaging of the type Iax SN 2012Z revealed a He-rich companion candidate (2014Natur.512...54M), and early spectra of some SNe Iax show He emission lines that likely originate in the circumstellar environment (2019MNRAS.487.2538J). However, reproducing the faintest SNe Iax within pure-deflagration models often requires very small ejecta masses, in tension with SN 2021fcg—the faintest Iax yet observed—whose inferred ejecta mass is comparable to that of brighter events (2021ApJ...921L...6K).
@@
An alternative channel is an ONe + CO WD merger in which only the CO component detonates, leaving the ONe WD as the remnant (2018ApJ...869..140K; Oskinova_et_al_20); this pathway generally does not reproduce light curves and spectra of SNe Iax as well as pure-deflagration models. Moreover, recently identified intermediate-luminosity events such as SN 2019muj (2021MNRAS.501.1078B; 2021PASJ...73.1295K; 2022ApJ...941...15M), SN 2022xlp (2025A&A...703A..64B), SN 2024pxl (2025arXiv250502943S; 2025ApJ...988..209H; 2025ApJ...989L..33K) and SN 2025qe (2025MNRAS.543.3731M) show features overlapping with those of both the bright and faint groups, suggesting a continuous character rather than two discrete subclasses. Taken together, these considerations keep the progenitor systems of SNe Iax under active discussion. Recent work further argues that in single-degenerate systems a deflagration–detonation transition may often fail, favoring Iax-like, low-energy outcomes over normal SNe Ia (2025arXiv250716907M). Clarifying the progenitors of SNe Iax is therefore central to the broader SN Ia progenitor problem.
@@
Figure 1: Decline rate $\Delta m_{15}$ versus peak absolute magnitude $M$ for normal SNe Ia and for the “bright” and “faint” subclasses of SNe Iax (reproduced from Fig. 11 of 2022MNRAS.511.2708S; filled/open symbols denote rr/$R$). The red region marks SN 1181; its values are not measured in $r/R$ but inferred from historical visual (naked-eye) estimates: a peak $M\simeq-12.5$ to 16-16 mag (2021ApJ...918L..33R; 2023MNRAS.523.3885S). The SN 1181 region is placed on the $r/R$ plane for comparison only.
@@
Several recently studied Type Iax SN remnants show late-time excess emission (e.g., SN 2012Z; 2022ApJ...925..138M). In some cases, their spectra/SEDs are better explained by a two-component model (e.g., SN 2014dt; 2016MNRAS.461..433F; SN 2012Z; 2025arXiv250401063S; and SN 2024pxl and SN 2024vjm; 2025ApJ...989L..33K). The two-component spectrum consists of a nebular component plus an additional blackbody-like continuum. The blackbody component is interpreted as arising from an optically thick wind launched from the bound remnant, as suggested for Pa 30. The wind could be powered by delayed radioactive decays of 56Ni and 56Co (2017ApJ...834..180S). Their characteristic velocities ($\sim 500$$1000~\mathrm{km~s^{-1}}$) are, however, far below the extreme wind inferred for the central WD in Pa 30. To account for this exceptionally fast wind, 2019ApJ...887...39K proposed a model involving a rapidly rotating, strongly magnetized WD that accelerates a carbon-ash outflow where the main energy source is shell carbon burning; the evolutionary pathway to such a configuration has not been systematically explored.
@@ 2 Methods
We investigate conditions for the ignition of fallback matter accreted onto the central WD inside Pa 30 to find the parameter space that allows for the delayed carbon ignition on timescales of $\sim 800$ yr. As discussed in Sec 1, the pure-deflagration model remains the leading progenitor scenario for SNe Iax, while merger channels involving an ONe WD and CO WD have also been proposed. Because simulations across these scenarios indicate that SN Iax ejecta are CO-rich (e.g., 2014MNRAS.438.1762F; 2018ApJ...869..140K), we restrict our analysis to accretion of matter containing only CO for simplicity. In pure-deflagration scenarios, the bound remnant at the center may be a CO or ONe WD, whereas in the WD merger channel the central remnant is generally an ONe WD. For Pa 30, prior estimates place the WD mass at $\sim 1.1$$1.3~M_{\odot}$ (e.g., 2019ApJ...887...39K; Lykou2022), which is more in line with typical masses of ONe WDs; CO WDs have masses around $\sim 0.6~M_{\odot}$ (1990sse..book.....K for a review.).
@@ 4.1 Implication on SN 1181 Progenitor system
If SN 1181 is a pure-deflagration SN Iax, a surviving He star companion is expected, however, none has been identified within Pa 30. Whether the surviving companion stays bound to the compact remnant depends on the amount of mass lost in the SN ejecta and the kick imparted to the remnant due to asymmetries in the explosion. Pure-deflagration simulations still leave the kick velocity of the remnant highly uncertain (e.g., 2012ApJ...761L..23J; 2014MNRAS.438.1762F), so it is not firmly known whether the WD–He star binary will be disrupted. However, because the ejecta mass is expected to be well below half of the total system mass, the system should, in most cases, remain bound. In the context of SNe Ia, it is known that when the WD in a WD-He star binary explodes, the He star companion can be shock-heated and subsequently expand; its radius may increase by a factor of $\sim 2$ on a timescale of $\sim 10^{6}$$10^{8}\,\mathrm{yr}$ (2019ApJ...887...68B); see also 2025ApJ...992..108W for the case of a He WD companion. Such shock heating and expansion of the He-star donor can make it easier for the companion to fill its Roche lobe, allowing Roche-lobe overflow to resume and mass transfer onto the bound WD remnant. Depending on the post-explosion orbital separation and the mass-transfer rate, the interaction may remain stable and produce a long-lived accreting binary, or become unstable, particularly if the WD remnant itself is thermally inflated, leading to a common-envelope like phase and ultimately a merger into a single object. Moreover, continued He accretion onto an ONe WD remnant may trigger unstable He burning, potentially producing recurrent He-shell flashes on the ONe WD sruface (2021RAA....21...34G). If the remnant can retain a significant fraction of the accreted material and grow in mass, it may eventually undergo accretion-induced collapse to a neutron star (e.g., 1991ApJ...367L..19N). Alternatively, if a sufficiently massive He layer accumulates on the surface, a He-shell detonation may occur, giving rise to a faint, rapidly evolving thermonuclear transient (i.e., an SN .Ia; 2007ApJ...662L..95B; 2010ApJ...715..767S).
If the binary system is not disrupted, two possibilities remain at least. First, if the WD is enshrouded by an optically thick wind with the photosphere at $\sim 10^{10}~\mathrm{cm}$, the He star could lie within it and be obscured. Second, the He star companion may lie outside the photosphere but be intrinsically faint—particularly at UV wavelengths—and has therefore escaped detection in existing observations. The two cases are examined in the following subsections.